random samples of size n=2 are drawn from a finite population consisting of the numbers 5,6,7,8,9. compute fot the mean, variance and standard deviation of the population, and the mean, variance and standard deviation of the mean


1
Expert's answer
2021-07-14T13:09:50-0400

The population values are "5, 6, 7, 8" "and" "9".

The sample size "n=2" .


"Population\:mean=\frac{5+6+7+8+9}{5}"

"Population\:mean=7"


"\:Variance=\frac{\left(5-7\right)^2+\left(6-7\right)^2+\left(7-7\right)^2+\left(8-7\right)^2+\left(9-7\right)^2}{5}"

"Variance=\frac{4+1+0+1+4}{5}"

"Variance=2"


"Standard\:deviation=\sqrt{2}" "=1.41"


The table for sampling distribution is shown below:




"The\:variance\:of\:sampling\:distribution=\frac{\left(5.5-7\right)+\left(6-7\right)+2\left(6.5-7\right)+2\left(7-7\right)+2\left(7.5-7\right)+\left(8-7\right)+\left(8.5-7\right)}{10}"

"Variance\:of\:sampling\:distribution\:is\:=\frac{2.25+1+0.5+0+0.5+1+2.25}{10}"

"Variance\:of\:sampling\:distribution\:is\:=0.75"


"Sample\:standard\:deviation=\sqrt{0.75}"

"Sample\:standard\:deviation=0.86"